# Validity of relativistic hydrodynamics beyond local equilibrium

### v1

@@ Abstract
In the 0+10+1D Bjorken scenario, we demonstrate that the exact evolution of non-equilibrium terms shares the same structural form as the gradient expansion, differing only through modified transport coefficients that reflect both initial data and free-streaming dynamics. Extending to 3+13+1D, we find that hydrodynamics remains effective not because the system is close to equilibrium, but because it interpolates smoothly between free streaming and collective behavior. This perspective offers a natural explanation for the remarkable success of hydrodynamics in modeling quark–gluon plasma evolution.
@@ I Introduction
Relativistic hydrodynamics is an effective theory that describes the macroscopic evolution of energy–momentum and conserved charges in a fluid at relativistic energies. It finds key applications ranging from astrophysical plasmas and cosmology to high-energy nuclear collisions.
The applicability of relativistic hydrodynamics to the quark–gluon plasma (QGP) is somewhat puzzling. In principle, hydrodynamics is expected to be valid only near local thermal equilibrium, in systems that are sufficiently large and slowly varying compared to microscopic scales. In contrast, the QGP produced in heavy-ion collisions is created in a highly anisotropic, far-from-equilibrium state. Despite these conditions, viscous hydrodynamics provides a quantitatively accurate description[1] of collective flow observables like particle spectra and flow coefficients [2, 3, 4, 5]. Even more surprisingly, flow-like correlations observed in small systems such as pPbp–Pb and high-multiplicity ppp–p [6, 7]collisions can also be reproduced within a hydrodynamic framework. These successes indicate that hydrodynamic behavior can emerge much earlier and in much smaller systems than expected. These observations raises the hydrodynamisation question: why does hydrodynamics work so well beyond its domain of validity?
@@ II Relativistic Hydrodynamics
$T^{\mu\nu}_{\text{eq}}=\epsilon u^{\mu}u^{\nu}-P\Delta^{\mu\nu}\,$, $n^{\mu}_{\text{eq}}=nu^{\mu}\,$, (1)
where $\epsilon$ is the fluid rest frame energy density, $P$ the rest frame fluid pressure, $n$ the number density and the four-velocity is normalized to $u^{\mu}u_{\mu}=1$. The tensor $\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}$ acts as a projection operator that maps any vector onto the subspace orthogonal to $u^{\mu}$. The conservation equations
@@
together with an equation of state of the form $\epsilon=\epsilon(n,P)$, provide a closed set of equations for describing the evolution of the system in local equilibrium. Alternatively, one may describe the equilibrium state using temperature ($T$) and chemical potential($\mu$) instead of $\epsilon$, $n$, and $P$.
@@ II.2 The Chapman-Enskog expansion
The Chapman-Enskog expansion assumes that the distribution function depends functionally on the fluid-dynamical variables—such as energy density $\epsilon$, number density $n$, anisotropy $\Pi^{\mu\nu}$ and their gradients- $f(\rho,\nabla\rho)$ or more generally, a hierarchy of moments $\rho_{n}$)—and that this dependence can be systematically expanded in powers of a small parameter $Kn\sim\lambda/L$,
@@ III.1 The Boltzmann Kernel
where $\xi_{0}=\int_{t_{0}}^{t}\frac{u\cdot p}{\tau_{R}}dt$’ is the damping factor. Notably, it mirrors the solution (III.1) under the identifications $v(p)\sim u\cdot p/\tau_{R}$, $\chi_{0}\sim\xi_{0}$ and $g\sim f_{\text{eq}}/\tau_{R}$. In fact any relaxation type model - $L[f]\to v(p)g[f]$ for the collision kernel will have a gradient expansion structurally similar to (III.1).
@@ III.2 Validity of the Gradient Expansion
Before discussing the validity of the relativistic gradient expansion, it is important to distinguish between the gradient expansion itself and relativistic hydrodynamics. The gradient expansion refers specifically to the asymptotic series obtained via a perturbative expansion in the Knudsen number ($Kn$). Relativistic hydrodynamics, on the other hand, encompasses not only this gradient series but also non-perturbative corrections—an aspect that will be made more rigorous in later sections. The expansions shown in (III.1) and (III.1) naturally include non-perturbative contributions—often labeled as “non-hydrodynamic” modes. This label, however, can be misleading as such modes are essential for ensuring both causality [42] and convergence [41] of the hydrodynamic theory. It is more accurate to interpret them as non-perturbative corrections, not non-hydrodynamic.
@@
The Chapman-Enskog gradient expansion remains valid as long as free-streaming effects are suppressed—i.e., when $\chi_{0}$ or $\xi_{0}\gg 1$—even when the gradients are large. We therefore argue that the applicability of the gradient expansion should not be judged by small gradients alone, but rather by how far ($\chi_{0}~\text{or}~\xi_{0}\gg 1$) it is from the free-streaming regime. In the simple 11D toy model this condition implies each other i.e $t/\delta>>1\implies\delta\frac{d}{dt}<<1$, but in general .
For a smooth, well-posed dynamical system, fixed points can only be approached asymptotically and never reached at finite time. The Chapman–Enskog gradient expansion, however, implicitly assumes that one can construct a solution valid at every finite time by expanding around the fixed point (the local equilibrium). This clashes with the actual dynamics: trying to self-consistently enforce the fixed point at finite time might be the origin of acausality.
@@ IV Hydrodynamics in $\mathbf{0}+1$} D
We consider a system with rotational and translational symmetry in the x–y plane and boost invariance along the z-axis. It is then convenient to work in the following Milne coordinate system,
@@
$T^{\mu\nu}=\begin{pmatrix}\epsilon&0&0&0\\ 0&P+\Pi+\frac{\pi}{2}&0&0\\ 0&0&P+\Pi+\frac{\pi}{2}&0\\ 0&0&0&P+\Pi-\pi\end{pmatrix}\,$, (35)
where $\Pi$ is the bulk pressure and the shear tensor is reduced to the form $\pi^{\mu\nu}=\text{diag}(0,\frac{\pi}{2},\frac{\pi}{2},\pi)$. The transverse pressure $P_{T}=P+\Pi+\frac{\pi}{2}$ and the longitudinal pressure is $P_{L}=P+\Pi-\pi$.
@@
$\partialderivative{f}{\tau}-\frac{p^{z}}{\tau}\frac{\partial f}{\partial p^{z}}=-\frac{1}{\tau_{R}}(f-f^{eq})\,.$ (36)
@@ IV.1 Moment Equations
Here, the index $n$ encodes the energy scaling, while $l$ characterizes the anisotropy in the momentum distribution. In particular, the number density ($n$) corresponds to the moment $\rho_{0,0}$, and the energy density ($\epsilon$) is given by $\rho_{1,0}$. The equilibrium moments $\rho_{n,l}^{eq}$ when $f$ has the Maxwell-Juttner form has the expression,
$\rho_{n,l}^{eq}(T,z)=\frac{T^{n+3}}{2\pi^{2}}\frac{G_{n+3,l}(z)}{2l+1}e^{\mu/T}\,$, (38)
@@ IV.2 Solution of the moment equations
Using the integrating factor method we can write down the solution (Appendix (B.1)),
@@
An exact expression for the free streaming solution can be obtained (see Appendix (B.1))
@@
The leading order late time behavior can be easily read off from the expression $g_{n,l}(\tau>>\tau_{0})\sim\left(\frac{\tau_{0}}{\tau}\right)^{2l+1}$, which is independent of $n$.
@@ IV.3 Gradient expansion of the moments
The system asymptotically approaches the gradient expansion, while the contribution from free-streaming modes—encoded in the second term—decays exponentially. The expression in brackets in (IV.3) quantifies the discrepancy between the exact initial condition and its gradient expansion. However, it is important to note that neither of the two terms in (IV.3) is convergent by themselves. This behavior mirrors the expansion obtained for the full distribution function, where the exact solution contains both a divergent asymptotic gradient series and a non-perturbative term carrying initial data.
@@ IV.4 Dynamics of anisotropy
$\rho_{n,l}=\rho_{n,l}^{eq}+\pi_{n,l}\,$ (52)
where $\rho_{n,l}^{eq}$ given in (38) is a function of $T$ and $\mu$. The evolution of $T$ and $\mu$ or alternatively $\epsilon$ and $n$ is obtained from the conservation equation with the Landau matching constraint,
$\rho_{0,0}$ $=\rho_{0,0}^{eq}(T,\mu)\,$, (53)
$\rho_{1,0}$ $=\rho_{1,0}^{eq}(T,\mu)\,.$
@@
$\mathbf{\hat{D}}\pi_{n,l}=-\frac{\pi_{n,l}}{\tau_{R}}-\mathbf{\hat{D}}\rho^{eq}_{n,l}(\tau)\,.$ (55)
@@
$\pi_{n,l}^{G}=\sum_{k=1}^{\infty}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{eq}_{n,l}(\tau)\,.$ (56)
@@
$\mathbf{\hat{D}}\pi_{n,l}^{G}$ $=-\frac{\pi_{n,l}^{G}}{\tau_{R}}-\mathbf{\hat{D}}\rho^{eq}_{n,l}$ (60)
@@
$\pi_{n,l}^{T}(\tau)\sim e^{-\xi_{0}}\sum_{n=0}^{\infty}\left(\sum_{l=0}^{k}\frac{\xi_{0}^{k}}{k!}\right)\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{n,l}(\tau)$ (62)
These observations are central to explaining the effectiveness of hydrodynamics even far from local equilibrium. They indicate that one can construct hydrodynamics starting from the gradient expansion, provided the transport coefficients are suitably adjusted. This approach yields evolution equations that reproduce the exact dynamics, and it closely mirrors the heuristic used in standard derivations of relativistic hydrodynamics. We demonstrate this explicitly by deriving the exact hydrodynamic equations in the following subsection.
@@ IV.5 Hydrodynamic Equations
Building on the insights from the previous section, we now derive the hydrodynamic evolution equations from the anisotropic moment equations. In the 0+10+1D setting, the relevant quantities are the energy density $\epsilon$, the bulk viscous pressure $\Pi$ and the shear pressure $\pi$. Our goal is to obtain a closed set of dynamical equations for these quantities.
@@
$\partial_{\tau}\epsilon=-\frac{\epsilon+P_{L}}{\tau}\,$, (63)
@@
$\partial_{\tau}\Pi+\frac{\Pi}{\tau_{R}}$ $=\frac{\Pi}{\tau}-\mathbf{\hat{D}}P-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}$ (64)
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}$ $=-(4-c_{s}^{2})\frac{\pi}{\tau}+2(2-c_{s}^{2})\frac{\Pi}{\tau}-\frac{\pi_{1,2}}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}$ (65)
where the operator $\mathbf{\hat{D}}P$ captures the pressure gradients and is given by,
$\mathbf{\hat{D}}P=-c_{s}^{2}\frac{(\epsilon+P+\Pi-\pi)}{\tau}-\frac{\kappa}{\tau}+\frac{3P}{\tau}-\frac{\rho^{eq}_{1,2}(\tau)}{\tau}\,$, (66)
with $c_{s}^{2}=\partialderivative{P}{\epsilon}$ representing the square of the speed of sound and $\kappa=n\partialderivative{P}{\text{n}}$ bulk modulus.
In deriving this expression, we used the trace of the energy-momentum tensor, which satisfies $T^{\mu}_{\mu}=\epsilon-3(P+\Pi)=m^{2}\rho_{-1,1}$. This relation serves as an effective equation of state. In the conformal limit ($m=0$), the bulk pressure vanishes ($\Pi=0$), and the equation of state reduces to $\epsilon=3P$.
The evolution equations (64) and (65) closely resemble those found in second-order hydrodynamics derived from traditional gradient expansion techniques [33, 34, 35, 36, 37, 38], though it’s important to stress that no approximation has yet been made in the derivation yet.
@@
To close the system of equations, the higher order moment couplings, $\pi_{-1,1}$ and $\pi_{1,2}$, must be expressed in terms of the hydrodynamic variables $\epsilon$,n, $\Pi$, and $\pi$. This can be achieved, in principle, by utilizing the expressions derived in equation (IV.3),
$\pi_{-1,1}$ $=e^{-\xi_{0}}e^{-\mathbf{\hat{K}}(\tau,\tau_{0})}\rho_{-1,1}(\tau_{0})-e^{-\xi_{0}}\rho_{-1,1}^{\text{eq}}(\tau)$
@@
$+\sum_{k=1}^{\infty}\frac{(-1)^{k}\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{-1,1}(\tau)\,$, (67)
$\pi_{~1,2}$ $=e^{-\xi_{0}}e^{-\mathbf{\hat{K}}(\tau,\tau_{0})}\rho_{1,2}(\tau_{0})-e^{-\xi_{0}}\rho_{1,2}^{\text{eq}}(\tau)$
$+\sum_{k=1}^{\infty}\frac{(-1)^{k}\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{1,2}(\tau)\,.$ (68)
However we see that a complete closure is not possible as the free streaming initial conditions $e^{-\xi_{0}}e^{-\mathbf{\hat{K}}(\tau,\tau_{0})}\rho_{-1,1}(\tau_{0})$ cannot be captured solely by gradient terms. Nonetheless, if
$\frac{e^{-\xi_{0}}\left(e^{-\mathbf{\hat{K}}(\tau,\tau_{0})}\rho_{n,l}(\tau_{0})-\rho^{\text{eq}}_{n,l}(\tau)\right)}{\pi_{n,l}}\ll 1$ (69)
for (n,l)=(1,1),(1,2)(n,l)=(-1,1),~(1,2), we can approximately close the system and write,
$\partial_{\tau}\Pi$ $+\frac{\Pi}{\tau_{R}}=\frac{\Pi}{\tau}-\mathbf{\hat{D}}P$
$+\frac{m^{2}}{3}\frac{1}{\tau}\left[\sum_{k=1}^{\infty}\frac{\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{-1,1}(\tau)\right]\,$, (70)
$\partial_{\tau}\pi$ $+\frac{\pi}{\tau_{R}}=-(4-c_{s}^{2})\frac{\pi}{\tau}+2(2-c_{s}^{2})\frac{\Pi}{\tau}$
$-\frac{1}{\tau}\left[\sum_{k=1}^{\infty}\frac{\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{1,2}(\tau)\right]$
$-\frac{m^{2}}{3}\frac{1}{\tau}\left[\sum_{k=1}^{\infty}\frac{\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{-1,1}(\tau)\right]\,.$ (71)
These expressions coincide with those obtained from the gradient expansion, with the crucial difference that the transport coefficients are rescaled by non-perturbative corrections arising from the $\gamma(k+1,\xi_{0})$ terms. Furthermore, the contributions $\left[\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{n,l}$ can always be expressed in terms of the hydrodynamic variables $T$, $\mu$, $\Pi$, and $\pi$ (see Appendix (B.3) for explicit formulas). This provides a rigorous understanding of the re-normalized transport coefficients proposed in previous studies [18, 21].
The key takeaway from the analysis is that
Hydrodynamic evolution $\sim\textit{gradient expansion}$
+ $+\textit{ modified transport coefficients }$
and the accuracy of this hydrodynamic approximation is quantified by the smallness of the ratio in (69).

### v2

# On the Validity of relativistic hydrodynamics beyond local equilibrium

@@ Abstract
In the 0+10+1D Bjorken scenario, we demonstrate that the exact evolution of non-equilibrium terms shares the same structural form as the gradient expansion, differing only through modified transport coefficients that reflect both initial data and free-streaming dynamics. Extending to 3+13+1D, we find that hydrodynamics remains effective not because the system is close to equilibrium, but because it interpolates smoothly between free streaming and collective behavior. This perspective offers a possible explanation for the remarkable success of hydrodynamics in modeling quark–gluon plasma evolution.
@@ I Introduction
Relativistic hydrodynamics is an effective theory that describes the macroscopic evolution of energy–momentum and conserved charges in a fluid at relativistic energies. It finds key applications ranging from astrophysical plasmas[1, 2, 3], relativity [4] and cosmology to high-energy nuclear collisions [5].
The applicability of relativistic hydrodynamics to the quark–gluon plasma (QGP) is somewhat puzzling. In principle, hydrodynamics is expected to be valid only near local thermal equilibrium, in systems that are sufficiently large and slowly varying compared to microscopic scales. In contrast, the QGP produced in heavy-ion collisions is created in a highly anisotropic, far-from-equilibrium state. Despite these conditions, viscous hydrodynamics provides a quantitatively accurate description[lare:1986qj] of collective flow observables like particle spectra and flow coefficients [6, 7, 8, 9]. Even more surprisingly, flow-like correlations observed in small systems such as pPbp–Pb and high-multiplicity ppp–p [10, 11]collisions can also be reproduced within a hydrodynamic framework. These successes indicate that hydrodynamic behavior can emerge much earlier and in much smaller systems than expected. These observations raises the hydrodynamisation question: why does hydrodynamics work so well beyond its domain of validity?
@@ II Relativistic Hydrodynamics
$T^{\mu\nu}_{\text{eq}}=\varepsilon u^{\mu}u^{\nu}-P\Delta^{\mu\nu}\,$, $n^{\mu}_{\text{eq}}=nu^{\mu}\,$, (1)
where $\varepsilon$ is the fluid rest frame energy density, $P$ the rest frame fluid pressure, $n$ the number density and the four-velocity is normalized to $u^{\mu}u_{\mu}=1$. The tensor $\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}$ acts as a projection operator that maps any vector onto the subspace orthogonal to $u^{\mu}$. The conservation equations
@@
together with an equation of state of the form $\varepsilon=\varepsilon(n,P)$, provide a closed set of equations for describing the evolution of the system in local equilibrium. Alternatively, one may describe the equilibrium state using temperature ($T$) and chemical potential($\mu$) instead of $\varepsilon$, $n$, and $P$.
@@ II.2 The Chapman-Enskog expansion
The Chapman-Enskog expansion assumes that the distribution function depends functionally on the fluid-dynamical variables—such as energy density $\varepsilon$, number density $n$, anisotropy $\Pi^{\mu\nu}$ and their gradients- $f(\rho,\nabla\rho)$ or more generally, a hierarchy of moments $\rho_{n}$)—and that this dependence can be systematically expanded in powers of a small parameter $Kn\sim\lambda/L$,
@@ III.1 The Boltzmann Kernel
where $\xi_{0}=\int_{t_{0}}^{t}\frac{u\cdot p}{\tau_{R}}dt$’ is the damping factor. Notably, it mirrors the solution (III.1) under the identifications $v(p)\sim u\cdot p/\tau_{R}$, $\chi_{0}\sim\xi_{0}$ and $g\sim f_{\text{eq}}$. In fact any relaxation type model - $L[f]\to v(p)g[f]$ for the collision kernel will have a gradient expansion structurally similar to (III.1).
@@ III.2 Validity of the Gradient Expansion
Before discussing the validity of the relativistic gradient expansion, it is important to distinguish between the gradient expansion itself and relativistic hydrodynamics. The gradient expansion refers specifically to the asymptotic series obtained via a perturbative expansion in the Knudsen number ($Kn$). Relativistic hydrodynamics, on the other hand, encompasses not only this gradient series but also non-perturbative corrections—an aspect that will be made more clear in later sections. The expansions shown in (III.1) and (III.1) naturally include non-perturbative contributions—often labeled as “non-hydrodynamic” modes. This label, however, can be misleading as such modes are essential for ensuring both causality [46] and convergence [45] of the hydrodynamic theory. It is more accurate to interpret them as non-perturbative corrections, not non-hydrodynamic.
@@
For a smooth, well-posed dynamical system, fixed points can only be approached asymptotically and never reached at finite time. The Chapman–Enskog gradient expansion, however, implicitly assumes that one can construct a solution valid at every finite time by expanding around the fixed point (the local equilibrium).
The Chapman-Enskog gradient expansion remains valid as long as free-streaming effects are suppressed—i.e., when $\chi_{0}$ or $\xi_{0}\gg 1$—even when the gradients are large. We therefore argue that the applicability of the gradient expansion should not be judged by small gradients alone, but rather by how far ($\chi_{0}~\text{or}~\xi_{0}\gg 1$) it is from the free-streaming regime. In the simple one-dimensional toy model these two conditions are equivalent, since $t/\delta\gg 1\implies\delta,\frac{d}{dt}\ll 1$, but this equivalence does not necessarily hold in more general cases.
@@ IV Hydrodynamics in $\mathbf{0}+1$} D
We consider a system with rotational and translational symmetry in the xy plane and boost invariance along the z-axis. It is then convenient to work in the following Milne coordinate system,
@@
$T^{\mu\nu}=\begin{pmatrix}\varepsilon&0&0&0\\ 0&P+\Pi+\frac{\pi}{2}&0&0\\ 0&0&P+\Pi+\frac{\pi}{2}&0\\ 0&0&0&P+\Pi-\pi\end{pmatrix}\,$, (35)
where $\varepsilon$ is the energy density, $P$ the equilibrium isotropic pressure, $\Pi$ is the bulk pressure and the shear tensor is reduced to the form $\pi^{\mu\nu}=\text{diag}(0,\frac{\pi}{2},\frac{\pi}{2},-\pi)$. The transverse pressure $P_{T}=P+\Pi+\frac{\pi}{2}$ and the longitudinal pressure is $P_{L}=P+\Pi-\pi$.
@@
$\partialderivative{f}{\tau}-\frac{p^{z}}{\tau}\frac{\partial f}{\partial p^{z}}=-\frac{1}{\tau_{R}}(f-f^{\rm eq})\,.$ (36)
@@ IV.1 Moment Equations
Here, the index $n$ encodes the energy scaling, while $l$ characterizes the anisotropy in the momentum distribution. In particular, the number density ($n$) corresponds to the moment $\rho_{0,0}$, and the energy density ($\varepsilon$) is given by $\rho_{1,0}$. The equilibrium moments $\rho_{n,l}^{\rm eq}$ when $f$ has the Maxwell-Juttner form has the expression,
$\rho_{n,l}^{\rm eq}(T,z)=\frac{T^{n+3}}{2\pi^{2}}\frac{G_{n+3,l}(z)}{2l+1}e^{\mu/T}\,$, (38)
@@ IV.2 Solution of the moment equations
Using the integrating factor method we can write down the solution (Appendix (B)),
@@
An exact expression for the free streaming solution can be obtained (see Appendix (B))
@@
with $(\cdot)^{(k)}$ being the rising Pochhammer symbol.
The leading order late time behavior can be easily read off from the expression $g_{n,l}(\tau>>\tau_{0})\sim\left(\frac{\tau_{0}}{\tau}\right)^{2l+1}$, which is independent of $n$. The analytic free streaming solution in (46) is compared with the numerical solution of the moment equations in Fig.1
Figure 1: The $\log-\log$ plot of the free streaming of the moment $\rho_{1,1}$ as a function of time for the analytical solution (dotted) obtained in (46) vs the numerical solution (solid) for the moment equation.
@@ IV.3 Gradient expansion of the moments
The system asymptotically approaches the gradient expansion, while the contribution from free-streaming modes encoded in the second term—decays exponentially. The expression in brackets in (IV.3) quantifies the discrepancy between the exact initial condition and its gradient expansion. However, it is important to note that neither of the two terms in (IV.3) is convergent by themselves. This behavior mirrors the expansion obtained for the full distribution function, where the exact solution contains both a divergent asymptotic gradient series and a non-perturbative term carrying initial data.
@@ IV.4 Dynamics of anisotropy
$\rho_{n,l}=\rho_{n,l}^{\rm eq}+\pi_{n,l}\,$ (52)
where $\rho_{n,l}^{\rm eq}$ given in (38) is a function of $T$ and $\mu$. The evolution of $T$ and $\mu$ or alternatively $\varepsilon$ and $n$ is obtained from the conservation equation with the Landau matching constraint,
$\rho_{0,0}$ $=\rho_{0,0}^{\rm eq}(T,\mu)\,$, (53)
$\rho_{1,0}$ $=\rho_{1,0}^{\rm eq}(T,\mu)\,.$
@@
$\partial_{\tau}\pi_{n,l}+\frac{\pi_{n,l}}{\tau_{R}}=-\mathbf{\hat{D}}\rho^{\rm eq}_{n,l}(\tau)-\mathbf{\hat{F}}\pi_{n,l}\,.$ (55)
@@
$\pi_{n,l}^{G}=\sum_{k=1}^{\infty}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\rm eq}_{n,l}(\tau)\,.$ (56)
@@
$\mathbf{\hat{D}}\pi_{n,l}^{G}$ $=-\frac{\pi_{n,l}^{G}}{\tau_{R}}-\mathbf{\hat{D}}\rho^{\rm eq}_{n,l}$ (60)
@@
$\pi_{n,l}^{T}(\tau)\sim e^{-\xi_{0}}\sum_{k=0}^{\infty}\left(\sum_{l=0}^{k}\frac{\xi_{0}^{k}}{k!}\right)\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{n,l}(\tau)$ (62)
These observations are central to explaining the effectiveness of hydrodynamics even far from local equilibrium. They indicate that one can construct hydrodynamics starting from the gradient expansion, provided the transport coefficients are suitably adjusted. This approach yields evolution equations that reproduce the exact dynamics, and it closely mirrors the heuristic used in standard derivations of relativistic hydrodynamics. We demonstrate this explicitly by deriving the exact hydrodynamic equations in the following subsection and approximating them with rescaled transport coefficients.
@@ IV.5 Hydrodynamic Equations
Building on the insights from the previous section, we now derive the hydrodynamic evolution equations from the anisotropic moment equations. In the 0+10+1D setting, the relevant quantities are the energy density $\varepsilon$, the bulk viscous pressure $\Pi$ and the shear pressure $\pi$. Our goal is to obtain a closed set of dynamical equations for these quantities.
@@
$\partial_{\tau}\varepsilon=-\frac{\varepsilon+P_{L}}{\tau}\,$, (63)
@@
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}$ $=-\frac{8}{3\tau}\pi+\frac{5}{3\tau}\Pi-\frac{\beta_{\pi}}{\tau}$
$-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}-\frac{\pi_{1,2}}{\tau}$ (64)
$\partial_{\tau}\Pi+\frac{\Pi}{\tau_{R}}$ $=-\left[\frac{4}{3}-c_{s}^{2}\right]\frac{\Pi}{\tau}+\left[\frac{1}{3}-c_{s}^{2}\right]\pi-\frac{\beta_{\Pi}}{\tau}$
$-\frac{m^{2}}{3}\frac{1}{\tau}\pi_{-1,1}\,.$ (65)
@@
The detailed derivation and the transport coefficients $\beta_{\pi}$ and $\beta_{\Pi}$ can be found in appendix C. The evolution equations (65) and (64) closely resemble those found in second-order hydrodynamics derived from traditional gradient expansion techniques [37, 38, 39, 40, 41, 42], prior to the implementation of moment closure. To close the system of equations, the higher order moment couplings, $\pi_{-1,1}$ and $\pi_{1,2}$, must be expressed in terms of the hydrodynamic variables $\varepsilon$,n, $\Pi$, and $\pi$. This can be achieved, in principle, by utilizing the expressions derived in equation (IV.3),
$\pi_{-1,1}$ $=e^{-\xi_{0}}\left(e^{-\mathbf{\hat{K}}(\tau,\tau_{0})}\rho_{-1,1}(\tau_{0})-\rho_{-1,1}^{\text{eq}}(\tau)\right)$
$+\sum_{k=1}^{\infty}\frac{\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{-1,1}(\tau)\,$, (66)
$\pi_{~1,2}$ $=e^{-\xi_{0}}\left(e^{-\mathbf{\hat{K}}(\tau,\tau_{0})}\rho_{1,2}(\tau_{0})-\rho_{1,2}^{\text{eq}}(\tau)\right)$
$+\sum_{k=1}^{\infty}\frac{\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{1,2}(\tau)\,.$ (67)
These expressions coincide with those obtained from the gradient expansion, with the crucial difference that each gradient term ( and therefore the transport coefficients) are rescaled by non-perturbative corrections arising from the $\gamma(k+1,\xi_{0})$ terms. The pure gradient terms in the expansion gives near equilibrium, long time (long wavelength) dynamics while the exponentially decaying corrections contains memory effects of early time (short wavelength) free streaming effects.
The gradient corrections due to the higher order moments enter the hydrodynamic evolution equations (65) and (64) is of the order of $\mathcal{O}(1/\tau^{k})$, $k\geq 2$ and follow the usual gradient expanded hydrodynamic structure with each order scaled by $\gamma(k+1,\xi_{0})$ factor. These are expressible in terms of the hydrodynamic variables $\varepsilon$, $n$, $\Pi$, and $\pi$ (see Appendix (C)). However the first term which includes the free streaming corrections expressed in equations (IV.5) is of the order of $\mathcal{O}(1/\tau)$. We can incorporate these terms into the transport coefficient $\beta_{\pi/\Pi}$.
$\beta_{\pi}$ $\to\beta_{\pi}~-(1/3)m^{2}\pi_{-1,1}^{(0)}-\pi_{1,2}^{(0)}\,$, (68)
$\beta_{\Pi}$ $\to\beta_{\Pi}-(1/3)m^{2}\pi_{-1,1}^{(0)}\,.$ (69)
We can get from (46)
$\pi_{-1,1}^{(0)}$ $\approx e^{-\xi_{0}}\left[~\sum_{k=0}^{\infty}\,K_{-1,1,k}(\tau,\tau_{0})\rho_{-1,1+k}(\tau_{0})-\rho^{\rm eq}_{-1,1}(\tau)\right]\,$, (70)
$\pi_{1,2}^{(0)}$ $\approx e^{-\xi_{0}}\left[~\sum_{k=0}^{\infty}\,K_{~1,2,k}(\tau,\tau_{0})\rho_{1,2+k}(\tau_{0})~-\rho^{\rm eq}_{~1,2}(\tau)\right]\,$, (71)
where the initial moments $\rho_{n,l+k}(\tau_{0})$ which contain the initial moment distribution, can be model parameters that can be obtained by fitting to data.
The gradient terms can in general can be written in the form,
@@
$\sum_{k=1}^{\infty}\frac{\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\rho^{\text{eq}}_{n,l}(\tau)$ $={\rm a}_{n,l}~\epsilon+{\rm b}_{n,l}~{\rm n}$
$+{\rm c}_{n,l}~\Pi+{\rm d}_{n,l}~\pi$ (72)
where ${\rm a}_{n,l},~{\rm b}_{n,l},~{\rm c}_{n,l},~{\rm d}_{n,l}$ are functions of the hydrodynamic variables $\epsilon,~n~,\Pi,\pi$. This will effectively close the equations (65), (64) and absorb the higher order contributions into the transport coefficients.